OpenStudy (itsbribro):

What is the standard form of the equation? 5x = 4y – 3 A. 5x – 4y + 3 = 0 B. 5x = 4y – 3 C. 5x + 3 = 4y D. 5x – 4y = -3

OpenStudy (anonymous):

A. Considering the Standard Form Equation is Ax+By+C=0

OpenStudy (itsbribro):

can you help me with some more if your not busy?

OpenStudy (anonymous):

Of course I would help you :)

OpenStudy (itsbribro):

What is the slope of the line 4x - 8y - 16 = 0 in simplest form?

OpenStudy (itsbribro):

i know its a fraction

OpenStudy (anonymous):

Ok so Id put it in this form of y=mx+b So It'll be -8y=-4x-16 Then divide it by -8 So it'll end up as y=1/2x+2. So your slope is 1/2

OpenStudy (itsbribro):

i was abt to type 1/2

OpenStudy (anonymous):

I mean it'll end up as y=1/2x-2 but its the slope that matters here so cx

OpenStudy (itsbribro):

yea thanks :)

OpenStudy (anonymous):

You're welcome. Anything else you need help with?

OpenStudy (itsbribro):

yea

OpenStudy (anonymous):

What is it you need help with? :)

OpenStudy (itsbribro):

What is the equation of the line expressed in slope-intercept form?

OpenStudy (itsbribro):

?????? @marie.serina09

OpenStudy (anonymous):

Ok I notice two points on the x and y axis: (0,-4) and (-2,0). Use these two points to get the slope. Once you have the slope, use one of the points and the slope in this equation:y=mx+b to find b. Then you would have your equation in slope-intercept form. And if Im incorrect, let me know. Im a bit fuzzy around this cx

OpenStudy (anonymous):

Your answer should be y=-2x-4

OpenStudy (itsbribro):

alright thx

OpenStudy (anonymous):

No problem :)

OpenStudy (itsbribro):

Mia deposits either $5 or $3.50 into her savings account. She made 14 deposits equaling a total of $59.50. How many times did she deposit $5 and how many times did she deposit $3.50?

OpenStudy (itsbribro):

Mia made____ deposits of $5 and ____ deposits of $3.50